OpenStudy (anonymous):

solve the system of equations by substitution 8y-1=x and 3x=2y

OpenStudy (anonymous):

plug the first equation in the second (since one variable is already isolated) 3(8y - 1) = 2y 24y - 2y = 3 22y = 3 y = 3/22 plug y in either equation, it should give the same answer (good way to check if you're right) x = 8(3/22) - 1 = 1/11 solution to system equation is (x, y) = (1/11, 3/22)